# LADICA: A Large Shared Display Interface for Generative AI Cognitive Assistance in Co-Located Team Collaboration

### v1

@@ 5.2. Apparatus
The study took place in-person within a usability lab. The participants were organized into small groups, each consisting of three or four individuals. LADICA ran on a 65-inch Phillips multi-touch display with an integrated Chromium browser. As discussed in section 4.4.1,each participant was also provided with an iPad to enable simultaneous content creation (especially text entry) and editing. All iPads and the shared touch display were logged into the same workspace in LADICA, ensuring content synchronization. This setup allowed for all actions performed on the large screen or on any individual iPad to be visible to all members of the group. Figure 12 illustrates the study apparatus setting.
@@ 6.1. Generative AI for Supporting Human-Human Collaboration
The design of LADICA demonstrates our human-centered approach that proritizes the facilitation and augmentation essential cognitive tasks in human collaboration rather than taking over cognitive tasks for teams in designing LLM-based assistants. For example, relational hints help teams recognize collaboration opportunities by promoting the development and progression of shared awareness. LADICA’s ability to extract crucial information and retrieve relevant elements from ongoing discussions enables teams to seamlessly align the working memory of the conversation with the external workspace. Automatic grouping through diverse user-focused affinity lenses enables teams to discern thematic patterns and comprehend the evolution of themes as collaboration progresses.
The study results of LADICA revealed opportunities of leveraging LLMs to support additional cognitive aspects in team collaboration. For instance, LLMs can be instrumental in detecting potential conflicts during team discussions or in highlighting areas of concern within the team’s proposed ideas, thereby encouraging an environment of critical analysis. Furthermore, LLMs can provide inspiration to each team member with regard to strategies to contribute to the ongoing discussion.

### v2

@@ 5.2. Apparatus
The study took place in-person within a usability lab. The participants were organized into small groups, each consisting of three or four individuals. LADICA ran on a 65-inch Phillips multi-touch display with an integrated Chromium browser. As discussed in section 4.4.1,each participant was also provided with an iPad to enable simultaneous content creation (especially text entry) and editing. All iPads and the shared touch display were logged into the same workspace in LADICA, ensuring content synchronization. This setup allowed for all actions performed on the large screen or on any individual iPad to be visible to all members of the group. We also set up an external video camera to record the study sessions. Figure 12 illustrates the study apparatus setting.
@@ 5.4.1. Post-study questionnaire
5.4.2. Interaction Log Analysis
We applied the interaction flow analysis method (avouris2003analysis) to analyze the interaction log data, conversation transcripts, and video recordings from study sessions, The goal was to understand how participants used LADICA during collaboration and to assess the fulfillment of the design goals. One of the authors reviewed the qualitative data to identify and categorize interaction patterns within the system.
Ideation and knowledge sharing
LADICA offers features such as group goal decomposition, query-based idea expansion, and relation-based idea expansion to encourage dynamic ideation and knowledge sharing (DG1). We observed that the participants used the group goal decomposition seven times throughout all study sessions. four of these instances occurred during the early discussion phase, where team members aimed to understand relevant topics or break down subtasks for the initially assigned topic (e.g., P4: “Let’s see what we need to consider for planning a spring break trip to Miami”). The remaining instances took place in the middle phase of collaboration, when participants focused on diving into topics of shared interest, such as “engaging students for in-class discussion” or “group budget management strategy”. After viewing the results of goal decomposition, participants tended to review the recommended subtasks sequentially, either exploring a specific topic collaboratively or delegating tasks to individual members.
Participants used query based idea expansion 18 times and relation based idea expansion 15 times throughout the sessions. When an individual triggered the idea expansion features, in 19 of the 33 instances, they sought input from other group members to decide which subtask to explore together. In the remaining cases, participants worked independently on creating notes under the subtasks and later brought them to the group discussion. We also observed that in three instances, participants dismissed the suggested subtasks and chose to list the subtasks themselves, as the generated suggestions were either too generic or included trivial details. This highlights the need for further fine-tuning the prompt strategies to guide the model toward an appropriate level of granularity.
Streamline discussion and workspace synchronization
LADICA offers discussion-based key information extraction and discussion-relevant idea retrieval to facilitate synchronization between discussion and workspace (DG2). We found that these two features were frequently used during the study sessions. All groups in the study activated the discussion-based key information extraction feature at the beginning of their task sessions. They used on this feature to automatically create notes for tracking key information, allowing them to focus on discussing the topic (e.g., P8: “we can use the system to help us track key ideas in our conversation”). However, participants observed that the system sometimes extracted insignificant information from their conversations, which they had to manually delete. This indicates the need for further refinement of the corresponding prompts and the addition of a filtering function.
All of the groups used discussion-relevant idea retrieval to keep track of relevant existing ideas while engaging in discussion. They examined the system-prompted ideas after having a conversation on a specific topic and identified connections between the existing ideas and the newly discussed points. Group 2 also alternated between using key information retrieval and extraction to efficiently synchronize discussion and workspace during their collaboration. This demonstrates the usefulness of these features in helping groups connect ongoing discussions with existing content.
Diverse analytical perspectives
The affinity lens layers offer various functions to enable diverse and comprehensive perspectives for analyzing created content (DG3). Participants used the multi-faceted affinity lens generation feature a total of seven times during the study sessions. They employed this feature to either identify common themes among their divergent ideas generated during individual brainstorming or to discover additional perspectives to explore based on existing ideas. After the affinity lens was generated, groups tended to focus on a small subset of lenses that the entire group was interested in and extended their discussion by enabling hierarchical affinity analysis on the sub-topics identified. We found that two groups used the idea snapshot rollback feature to return to a previous state, as they found that the discussion had become too narrow or had diverged from the main topics they were assigned. We also observed that Group 1 did not use the affinity lens feature after trying it once because members felt that identifying common themes on their own would provide an opportunity for collaboration and help develop mutual understanding.’
Collaborative cues
LADICA fosters group discussion by offering relation hints among ideas and group-based ideation prompts (DG4). During our study, participants used relation hints 10 times to gain insight into the relationships between different members’ contributions. We observed that highlighting these relationships motivated group members to discuss the connected ideas and collaborate to expand on them together. However, since the system could present an overwhelming number of relations among ideas created by multiple members, some participants chose not to use this feature again, as they felt it might disrupt others’ ongoing work and clutter the interface. This suggests the need to refine the number of hints provided at a time. Participants activated group-based ideation hints 12 times throughout the study sessions, primarily using the feature when they identified topics of common interest and sought prompts for further discussion. In five instances, participants directly included the suggested ideas, while in the remaining seven cases, they used the hints as inspiration to create their own ideas.
@@ 6.1. Generative AI for Supporting Human-Human Collaboration
The design of LADICA demonstrates our human-centered approach that prioritizes the facilitation and augmentation essential cognitive tasks in human collaboration rather than taking over cognitive tasks for teams in designing LLM-based assistants. Our three-layer conceptual framework (Figure 2) echoed He et al.’s opinion (he2024ai) that AI should play dynamic roles in facilitating human collaborations. Unlike many prior work that proactively recommend ideas for inspiration, LADICA preserves user initiatives to generating ideas, while providing scaffolding assistance for idea expansion, articulation and task structure decomposition. The participants found that these forms of assistance helped to bootstrap the initial ideation phase of the collaboration. However, they noted that with static prompting templates, the generative ideation support was sometimes off-topic, biased towards certain topics or overly generic. This suggests a need to incorporate real-time contextual group-generated information to enhance the relevance of the scaffolding prompts.
In addition to serving as an ideation stimulus, LADICA also leverages a generative AI model to support collective analytical and reflective thinking based on the shared workspace. Our study found that by facilitating idea grouping, connection, and retrieval during collaboration, the generative AI model helps teams build a shared mental model, align internal and external memory, and initiate cooperation. However, participants also expressed concerns that the potential over-reliance on AI support could hinder the natural discovery of collectively interesting topics through iterative discussion and occasionally reduce group engagement (P5, P9), particularly since the current system lacks detailed rationale or comparisons of elements within the same group to justify the suggested grouping. Besides, while participants appreciated the relational hints for highlighting discussion opportunities, some (P2, P6, P9) felt that the system often generated too many hints, which increased their mental load and distracted them during individual brainstorming. These challenges highlight the need for refined prompt strategies and design efforts to enhance user control and explainability in analytical processes and determine user preferences for the number of suggestions.
@@ 8. Conclusion
Acknowledgements.
This work was supported in part by NSF grants CNS-2426395, DRL-2437113, CCF-2211428, an NVIDIA Academic Hardware Grant, a Google Research Scholar Award, a Notre Dame-IBM Technology Ethics Lab Award, and a University of Notre Dame Strategic Framework Teaching Grant. Any opinions, findings or recommendations expressed here are those of the authors and do not necessarily reflect views of the sponsors. We would like to thank Xu Wang, Ron Metoyer and Diego Gómez-Zará for useful discussions.